### v1

# PGA: Personalizing Grasping Agents with Single Human-Robot Interaction

Abstract
Language-Conditioned Robotic Grasping (LCRG) aims to develop robots that ground and grasp objects based on natural language instructions. While robots capable of recognizing personal objects like “my wallet” can interact more naturally with non-expert users, current LCRG systems primarily limit robots to understanding only generic expressions. To this end, we introduce a task scenario GraspMine with a novel dataset that aims to locate and grasp personal objects given personal indicators via learning from a single human-robot interaction. To address GraspMine, we propose Personalized Grasping Agent (PGA), that learns personal objects by propagating user-given information through a Reminiscence—a collection of raw images from the user’s environment. Specifically, PGA acquires personal object information by a user presenting a personal object with its associated indicator, followed by PGA inspecting the object by rotating it. Based on the acquired information, PGA pseudo-labels objects in the Reminiscence by our proposed label propagation algorithm. Harnessing the information acquired from the interactions and the pseudo-labeled objects in the Reminiscence, PGA adapts the object grounding model to grasp personal objects. Experiments on GraspMine show that PGA significantly outperforms baseline methods both in offline and online settings, signifying its effectiveness and personalization applicability on real-world scenarios. Finally, qualitative analysis shows the effectiveness of PGA through a detailed investigation of results in each phase.
I INTRODUCTION
Empowering robots with the ability to comprehend human natural language presents a formidable yet vital challenge within the realms of AI and Robotics [1]. This capability, which involves understanding and executing instructions couched in human language, allows for more intuitive human-robot interactions. Addressing this capability, the field has extensively studied Language-Conditioned Robotic Grasping (LCRG) [2, 3, 4], which focuses on robotic systems that ground and grasp objects based on language instructions.
Current approaches in LCRG predominantly rely on generic expressions when describing objects for manipulation, leading to less intuitive human-robot communication. For instance, a person might instinctively instruct, “Get my AirPods”. However, with current robotic systems that operate on generic understandings, one might need to instruct, “Get the white pebble-shaped object in front of the blue cup”. This forces users to craft their instructions to suit the robot’s limited knowledge base, what can be termed as robot-centric directives. However, many non-expert users find this robot-centric directives both unfamiliar and cumbersome. To truly facilitate more intuitive user-centric interactions, robots should have more shared personal knowledge with their users beyond the generic, understanding personalized expressions.
To address this problem, we introduce a novel personalized task scenario of LCRG called GraspMine with a benchmark dataset. GraspMine aims to locate and grasp personal objects given a personal indicator, e.g., “my sleeping pills”, that robotic systems with generic knowledge may not handle properly. According to the field of Personalization [5, 6], personalizing with a few examples is a common scenario since collecting a lot of personalized data for each user is infeasible in real-world applications. Considering this aspect, GraspMine requires robots to learn personal objects via minimal human-robot interaction, i.e., a one-time verbal introduction of a personal object, as depicted in Fig. 1. In GraspMine dataset, we consider learning of 96 personal objects, each with its own indicator. GraspMine contains five distinct test splits, each sample including an image containing multiple objects, a personal indicator, and the associated object’s coordinates.
Figure 1: Personalizing grasping agents with single human-robot interaction. Upon the user’s introduction of a personal object, it retrieves the identical objects from its visual reminiscence. Leveraging the retrieved objects, the robot subsequently engages in an integrated learning process of the personal object. The personalized grasping agent can finally comprehend the personalized expression, grasping the personal object.
To address GraspMine, we propose Personalized Grasping Agent (PGA), that learns personal objects by propagating the information acquired from a user through a Reminiscence. Specifically, PGA first constructs the Reminiscence—a collection of raw images from the user’s environment. Then PGA acquires personal object information in two successive steps: human-robot interaction and robot-object interaction. In human-robot interaction, a user introduces a personal item to the robot with a personal indicator, e.g.,my sleeping pills”. In robot-object interaction, the robot inspects the object from multiple perspectives, obtaining diverse images of personal objects. Leveraging the acquired information of the personal object, PGA pseudo-labels objects in the Reminiscence by propagating personal indicators based on the visual features of the objects, which stems from label propagation algorithm [7]. We term this Propagation through Reminiscence. Harnessing the data obtained from the interactions and the pseudo-labeled objects in Reminiscence, PGA adapts the object grounding model, and with the personalized model, it adeptly grasps the user-specified personal objects.
To validate the effectiveness of PGA, we conduct both offline (i.e., object discovery) and online (i.e., real-world LCRG) experiments on GraspMine. In offline experiments, PGA outperforms several baseline methods, and even shows comparable results with a fully-supervised method that employs extensive annotations within the Reminiscence. Moreover, PGA also robustly performs in an online setting, demonstrating its real-world personalization applicability. Concluding our analysis qualitatively, we present visual examples from various phases of PGA, offering insights into the mechanisms that contribute to its robust performance.
II RELATED WORK
Language-Conditioned Robotic Grasping (LCRG) focus on the robot’s capability to grasp objects based on human instructions given in natural language. Numerous studies [8, 9, 4, 10, 11, 12, 3, 13] have employed fully supervised training approaches to teach robots how to locate objects based on natural language instructions, making use of public datasets [14, 15, 16]. However, these supervised learning methods reach their limit when robots have to recognize or interact with personal objects absent during the training. A study [17] curated its own dataset comprising unique objects. Still, it depended on a supervised learning, which necessitates exhaustive annotations of language instructions. GVCCI [2] introduced a paradigm shift by unveiling an unsupervised approach to adapt to the user’s personal objects. While GVCCI enabled robots to automatically learn the visual nuances of personal objects, it remained constrained. The learning was heavily reliant on basic-level categories [18] and attributes inferred by pretrained classifiers [19, 20]. A major limitation of such methodologies is their grounding of visual objects in basic-level categories. This forces non-expert users to use robot-centric instructions, a pain point discussed in Sec. I. Some studies [21, 22] delved deeper, augmenting robots’ knowledge with semantic or intended meanings of objects, typified by phrases like “I am thirsty”. While these studies enable robots to grasp user intentions, they also rely on fully-supervised learning, necessitating extensive annotations. Unlike previous works, PGA can efficiently learn the user’s personal objects and their corresponding language descriptors with just a single user-robot interaction.
Personalization has become an important factor within various domains of Machine Learning [23, 24, 25, 26, 27, 28, 29, 30]. Personalization has also been studied in the field of Robotics, such as dressing assistance [31, 32, 33] and tidy-up task [34]. While these works of Personalization focus on the user’s personal preferences, our work focus on transferring the user’s personal knowledge, standing as a pioneering work in the realm of personalized LCRG. Two studies [5, 6] align closely with our research objectives, focusing on learning visual and linguistic representations of personal objects. The work [6] utilizes personal videos paired with transcripts to learn personal objects and their corresponding names, without the need for explicit human labeling. Yet, their focus lies on temporal grounding in videos, overlooking the crucial aspect of spatial grounding. This limitation renders their approach unsuitable for LCRG, where pinpointing the precise object location is imperative. The work [5] broadens this focus to include spatial grounding. However, they assume that fine-grained annotations (e.g., segmentation masks) and personal object names are readily available. In contrast, PGA only requires a single human-robot interaction without any further annotations. So, non-expert users can easily be involved in collecting the training data due to the intuitive and streamlined interface.
III METHOD
Figure 2: Overview of Personalized Grasping Agent (PGA). (a) Initially, PGA gathers a collection of raw images, termed as Reminiscence, from the user’s environment. With the personal indicator provided by the user from (b), unlabelled objects in the Reminiscence are pseudo-labelled via (c) Propagation through Reminiscence. It’s vital to note that certain objects, particularly those not introduced by the user (e.g., non-personal objects), remain unlabelled by our algorithm. Ultimately, PGA employs all the object nodes with labels (colored nodes) to train the Personalized Object Grounding Model.
III-A Reminiscence Construction
Personalized Grasping Agent (PGA) initiates by constructing the Reminiscence, as illustrated in Fig. 2-(a). PGA first collects a set of $N$ images from the user’s environment, represented as =$\mathcal{R}=\{I^{\mathcal{R}}_{n}\}_{n=1}^{N}$, which we term as Reminiscence. PGA detects all the objects $\{o^{\mathcal{R}}_{m}\}_{m=1}^{M}$ from $\mathcal{R}$, via an off-the-shelf object detector [20], and constructs a node for each detected objects, embedding them with the pretrained visual encoder $f(\cdot)$, DINO [35]. These object nodes, represented as $O^{\mathcal{R}}=\{\mathbf{o}^{\mathcal{R}}_{m}\}_{m=1}^{M}=\{f(o^{\mathcal{R}}_{m})\}_{m=1}^{M}$ where $M\gg N$, are tagged with the meta-information – an image from $\mathcal{R}$ and the bounding box (bbox) coordinates of the detected object – which will be leveraged in subsequent processes. Note that nodes in $O^{\mathcal{R}}$ are without the object’s personal indicators.
III-B Object Information Acquisition
To acquire the information of personal objects, PGA goes through two successive steps as shown in Fig. 2-(b): human-robot interaction and robot-object interaction. In human-robot interaction, the user introduces their personal object $o$ by displaying the object and verbally describing it. The description consists of two indicators: a general indicator $u^{G}_{o}$, e.g., “the object in front”, and a personal indicator $u_{o}$, e.g., “my sleeping pills”. Using its visual perception $\mathcal{I}_{o}$ and $u^{G}_{o}$, PGA predicts the bbox coordinates $b_{o}$ of the object via GVCCI [2]. PGA then constructs an object node $\mathbf{o}=f(o)$ using pretrained visual encoder $f(\cdot)$ [35]. This node is tagged with the meta-information including $\mathcal{I}_{o}$, $b_{o}$, and $u_{o}$.
In robot-object interaction, PGA subsequently examines the personal object $o$ from multiple views, aiming to capture its distinctive visual appearance. By leveraging $b_{o}$, PGA grasps and rotates the object capturing $A$ images $\{\mathcal{I}_{o,a}\}_{a=1}^{A}$ for each object, each image offering a distinct view of the object (see the bottom right of Fig. 2). As in the earlier step, PGA predicts an object bbox coordinates, $\{b_{o,a}\}_{a=1}^{A}$ of a set of objects $\{{o_{a}}\}_{a=1}^{A}$. PGA constructs a set of object nodes $O^{view}=\{{\mathbf{o}_{a}}\}_{a=1}^{A}=\{{f(o_{a})}\}_{a=1}^{A}$, where each node $\mathbf{o}_{a}$ is tagged with the meta-information $\mathcal{I}_{o,a}$, $b_{o,a}$, and $u_{o}$. Finally, each personal object is associated with $\mathbf{o}$ and $O^{view}$, i.e., $A+1$ nodes that are tagged with personal indicators. Across all personal objects, the aggregation of nodes with personal indicators forms $O^{\mathcal{K}}$, while collective personal indicators form $U$, as described in lines 1-6 of Algorithm 1.
III-C Propagation through Reminiscence
Propagation through Reminiscence draws inspiration from the semi-supervised method of label propagation [7, 36], as detailed in Algorithm 1 and depicted in Fig. 2-(c). The goal is to pseudo-label (tag) the missing indicators in $O^{\mathcal{R}}$ by propagating personal indicators from $O^{\mathcal{K}}$. Specifically, for every node $\mathbf{o}^{\mathcal{R}}\in O^{\mathcal{R}}$, PGA computes an affinity score, $S(\mathbf{o}^{\mathcal{R}},u)$, for each possible personal indicator $u\in U$. This score is the average cosine similarity $\phi$ between node $\mathbf{o}^{\mathcal{R}}$ and every known object node $\mathbf{o}^{\mathcal{K}}$ tagged with indicator $u$;
$\phi(\mathbf{o}^{\mathcal{K}},\mathbf{o}^{\mathcal{R}})=\frac{\mathbf{o}^{\mathcal{K}}\cdot\mathbf{o}^{\mathcal{R}}}{\|\mathbf{o}^{\mathcal{K}}\|_{2}\times\|\mathbf{o}^{\mathcal{R}}\|_{2}}=\frac{f(o^{\mathcal{K}})\cdot f(o^{\mathcal{R}})}{\|f(o^{\mathcal{K}})\|_{2}\times\|f(o^{\mathcal{R}})\|_{2}};$ (1)
$S(\mathbf{o}^{\mathcal{R}},u)=\frac{1}{N(u)}\sum_{\mathbf{o}^{\mathcal{K}}\in O^{\mathcal{K}}}I(u_{\mathbf{o}^{\mathcal{K}}}=u)\cdot\phi(\mathbf{o}^{\mathcal{R}},\mathbf{o}^{\mathcal{K}});$ (2)
where $N(u)$ denotes the number of $\mathbf{o}^{\mathcal{K}}$s tagged with indicator $u$ and $I(\cdot)$ is the indicator function that takes the value of 1 if the condition is true, and 0 otherwise. Note that $u_{\mathbf{o}^{\mathcal{K}}},u\in U$. Based on these scores, PGA assigns the indicator $u$ with the highest affinity score to $\mathbf{o}^{\mathcal{R}}$; $u_{o^{\mathcal{R}}}={\text{argmax}}\{S(\mathbf{o}^{\mathcal{R}},u)~{}|~{}u\in U\}$; if the highest score is above the certain threshold. After pseudo-labeling $\mathbf{o}^{\mathcal{R}}$ with the most likely personal indicator, $O^{\mathcal{K}}$ is updated to incorporate the propagated node $\mathbf{o}^{\mathcal{R}}$, and this updating process is iteratively performed until the proportion of altered nodes is less than 10% – until saturated. Propagation through Reminiscence allows PGA to effectively gather pseudo-labeled examples, thereby fostering robust, data-rich learning of the user’s personal object.
Algorithm 1 Propagation through Reminiscence
1:The set of user’s personalized objects $\mathcal{O}$
2:Visual Encoder $f(\cdot)$ [35]
3:Nodes $O^{\mathcal{R}}\leftarrow\{\mathbf{o}^{\mathcal{R}}_{m}\}_{m=1}^{M}$ from $\mathcal{R}$triangleright (Sec. III-A)
4:Set of personal indicators $U=\emptyset$
5:Nodes with indicators $O^{\mathcal{K}}=\emptyset$
6:for $o\leftarrow\mathcal{O}$ do $\triangleright$ (Sec. III-B)
7:     $\mathbf{o}\leftarrow f(o)$ and $u_{o}$ from human-robot interaction
8:     Get $\{{o_{a}}\}_{a=1}^{A}$ from robot-object interaction
9:     $O^{view}\leftarrow\{{f(o_{a})}\}_{a=1}^{A}$ tagged with $u_{o}$
10:     $O^{\mathcal{K}}\leftarrow O^{\mathcal{K}}\cup\{\mathbf{o}\}\cup O^{view}$ and $U\leftarrow U\cup\{u_{o}\}$
11:end for
12:repeatnormal-$\triangleright$ (Sec. III-C)
13:     for $\mathbf{o}^{\mathcal{R}}\leftarrow O^{\mathcal{R}}$ do
14:         $\forall u\in U$, set $S_{u}=0$, $N_{u}=0$
15:         for $\mathbf{o}^{\mathcal{K}}\leftarrow O^{\mathcal{K}}$ do
16:              $u\leftarrow u_{\mathbf{o}^{\mathcal{K}}}$
17:              $S_{u}\leftarrow S_{u}+\phi(\mathbf{o}^{\mathcal{K}},\mathbf{o}^{\mathcal{R}})$ and $N_{u}\leftarrow N_{u}+1$
18:         end for
19:         $\forall u\in U$, $S(\mathbf{o}^{\mathcal{R}},u)\leftarrow S_{u}/N_{u}$
20:         if max $\{S(\mathbf{o}^{\mathcal{R}},u)~{}|~{}u\in U\}$ $>$ threshold then
21:              $u_{o^{\mathcal{R}}}\leftarrow\text{argmax}\{S(\mathbf{o}^{\mathcal{R}},u)~{}|~{}u\in U\}$
22:              $O^{\mathcal{K}}\leftarrow O^{\mathcal{K}}\cup\{\mathbf{o}^{\mathcal{R}}\}$
23:         end if
24:     end for
25:until Ratio of changed indicator of node $<$ 10%
III-D Personalized Object Grounding Model
The Personalized Object Grounding Model takes an image and a natural language indicator to infer the location of the personal object. We adopt the architecture of OFA [37], a Transformer [38] based Vision-Language model. The model’s optimization occurs by minimizing
$\mathcal{L}=-\sum_{\mathbf{o}^{\mathcal{K}}\in O^{\mathcal{K}}}logP_{\theta}(b_{\mathbf{o}^{\mathcal{K}}}|\mathcal{I}_{\mathbf{o}^{\mathcal{K}}},u_{\mathbf{o}^{\mathcal{K}}}),$ (3)
the negative log-likelihood of the bounding box for every object node $\mathbf{o}^{\mathcal{K}}\in O^{\mathcal{K}}$. Note that $O^{\mathcal{K}}$ is not from Sec. III-B but after Propagation through Reminiscence (Sec. III-C).
III-E Personalized Object Grasping
The process of Personalized Object Grasping is initiated when a user instructs the robot to grasp a personal object as seen in Fig. 1. To execute this, PGA first infers the 2D coordinates of the queried object using the Personalized Object Grounding Model. Upon obtaining these 2D bounding box coordinates, PGA calculates the 3D segmented object coordinates by leveraging point cloud data and the RANSAC [39] algorithm. Specifically, it translates the 2D bounding box into a 3D spatial configuration using the point cloud, followed by segmenting the points of the object within the 3D bounding box via RANSAC. Lastly, PGA grasps the object by computing a trajectory of the robot arm [40].
IV EXPERIMENTS
IV-A Dataset
Our proposed dataset, GraspMine, is curated with 96 personal objects and 100+ everyday objects, encompassing Reminiscence images =$\mathcal{R}=\{I_{n}^{\mathcal{R}}\}_{n=1}^{N}$ and five distinct Test splits.
Reminiscence $\mathcal{R}$ involves $N=400$ raw images, each containing multiple objects. We annotated objects in the Reminiscence for analyzing the propagation ability of models and for training the Supervised method in Sec. IV-C and did not use the annotation in any process in training.
Test set is categorized into five distinct splits as follows. Each data comprises an image, a personal indicator, and its corresponding object bounding box (bbox) annotation.
  • Heterogeneous split incorporates scenes with randomly selected objects as exemplified in Fig.4-A. It consists of 60 images with 120 personal indicators and bboxes.
  • Homogeneous split incorporates scenes with similar-looking objects with the same category, making discrimination more challenging as exemplified in Fig.4-B,C. It consists of 60 images with 120 indicators and bboxes.
  • Cluttered split contains 106 images containing highly cluttered objects as exemplified in Fig.4-D,E, and a single personal indicator and bbox per image. Images of this split originate from the IM-Dial dataset [22].
  • Paraphrased split comprises all the images of Heterogeneous, Homogeneous, and Cluttered split with each personal indicator paraphrased by the annotators.
  • Generic split is sourced from the VGPI dataset presented in GVCCI [2]. This dataset features images comprising generic objects, each annotated with indicators based on their basic-level object categories. Within the VGPI dataset, we specifically utilized the Test-E split, which aligns with our environmental settings.
IV-B Compared Methods
OFA [37] is a state-of-the-art visual grounding model pre-trained using a public dataset [14]. This model has never been exposed to the user’s personal objects, positioning it as the most basic compared method. It primarily relies on generic cues (e.g., “tennis ball”) present in personal indicators (e.g., “my Djokovic tennis ball”) to make predictions.
GVCCI [2] automatically generates generic expressions of the objects in the Reminiscence and harnesses generated data for training. It shares the same model architecture with OFA.
Direct model is trained directly with the nodes $\mathbf{o}$ from the human-robot interaction in Sec. III-B. This model does not employ robot-object interaction in Sec. III-B and the Propagation through Reminiscence (Sec. III-C), resulting in training the model with a single sample per personal object.
Passive model is a variant of PGA, excluding robot-object interaction in Sec. III-B. In other words, Propagation through Reminiscence (Sec. III-C) is conducted only with the nodes from human-robot interaction in Sec. III-B.
Supervised model is trained with a fully-supervised approach that utilizes the ground-truth annotations (i.e., object location coordinates and personal indicators) in Reminiscence. It is regarded as the model showing the upper bound performance in the following experiments.
TABLE I: Results of offline experiments. Results show the personal object grounding score in GraspMine dataset. Bold scores represent the highest performance, while underlined scores the second-best results. The ‘interaction’ column represents the average number of interactions per object that users had with the robot to teach personal objects. ‘annotated’ indicates the total human annotations used in each method, while ‘utilized’ specifies the number of triplets - image, object bounding box, and personal indicator - employed for training the models. It’s noteworthy that subtracting the ‘annotated’ count from the ‘utilized’ count reveals the amount of triplet generated by each method.
Heterogeneous Homogeneous Paraphrased Cluttered Generic
Method interaction annotated utilized $\text{\footnotesize IoU}_{>0.5}$ $\text{\footnotesize IoU}_{>0.8}$ $\text{\footnotesize IoU}_{>0.5}$ $\text{\footnotesize IoU}_{>0.8}$ $\text{\footnotesize IoU}_{>0.5}$ $\text{\footnotesize IoU}_{>0.8}$ $\text{\footnotesize IoU}_{>0.8}$ $\text{\footnotesize IoU}_{>0.5}$
OFA [37] - - - 49.2 47.5 23.7 20.3 35.8 34.1 34.6 65.7
GVCCI [2] - - - 59.3 55.1 30.5 23.8 42.3 38.9 44.3 79.1
Direct 1.1 96 96 60.2 55.1 37.3 27.1 48.3 42.5 46.4 -
Passive 1.1 96 4,828 89.0 76.3 68.6 53.4 73.6 62.7 62.4 -
PGA (ours) 1.1 96 6,492 91.5 81.4 70.3 61.9 74.7 68.2 64.5 79.1
Supervised 91.3 8,763 8,763 97.5 90.7 92.4 83.1 84.8 79.0 71.4 -
IV-C Offline Experiment
Evaluation Protocol. In the offline experiment, we study the PGA’s proficiency in grounding the target object given a personal indicator. By following standard practice [2, 37], we assess the Personalized Object Grounding Model using the Intersection over Union (IoU) score, a key visual grounding metric, which calculates the overlap between the predicted and ground truth bounding boxes. While we present the percentage of predictions surpassing an IoU of 0.5, we also emphasize IoU above 0.8 to ensure a stricter alignment between the predicted and actual regions, acknowledging the precision needed for successful object grasping. For the Cluttered split, we only reported scores with an IoU exceeding 0.8 since boxes with an IoU above 0.5 frequently included multiple objects, rendering the score less meaningful.
Results. The offline results, presented in Tab. I, demonstrate that PGA substantially outperforms methods reliant on generic object knowledge, i.e., OFA and GVCCI. PGA’s advantage is also evident when compared to the Direct model, which solely uses object information acquired from human-robot interaction and ignores data from its Reminiscence. The improvement over the Passive model underscores the impact of robot-object interaction, emphasizing the importance of capturing objects from different views. Remarkably, even when compared to the Supervised method, which requires extensive annotations (91.3 annotations of per object on average) of Reminiscence, PGA, which only requires 1.1 interaction per object on average, shows comparable results. We also assess PGA in grounding objects when queried with generic instructions, using the Generic split. Remarkably, even after acquiring personal knowledge, PGA’s ability to ground objects through generic instructions remains impeccably intact. In short, by utilizing Reminiscence and the robot’s manipulative power, PGA proficiently grounds personal objects with learning from a single human-robot interaction, outperforming the baseline methods and showing comparable performance to the Supervised approach, all while retaining its knowledge of generic expressions.
Figure 3: Impact of Reminiscence size. PGA’s offline scores based on the number of raw images utilized from Reminiscence, from ‘0’ to ‘400’.
Impact of Reminiscence size. Shown in Fig. 3, our exploration delves into the relationship between PGA’s performance and the size of the Reminiscence — the number of raw images utilized. A size of ‘0’ is equivalent to the Direct model. We experimented on a log scale, examining sizes of 25, 100, and 400, with each configuration undergoing three separate trials. Our findings suggest a clear pattern: as PGA is exposed to an increasing number of scenes from the user’s environment, its performance is incrementally enhanced. This indicates that, as the robot collects more raw images from the environment, its performance is expected to be improved without necessitating further human supervision.
Analysis of Propagation through Noisy Reminiscence. Detecteed objects in the Reminiscence are noisy as they are predicted using an off-the-shelf object detector [20]. Therefore, in this analysis, we investigate how many noisy objects are pseudo-labelled with Propagation through Reminiscence. We divided noisy objects into two categories: 1) ambiguous boxes, i.e., boxes that only include a part of an object (e.g., Fig. 4-(a)) or boxes that surround multiple objects, and 2) invalid boxes, i.e., non-objects such as line stickers on desks or a robot arm (e.g., Fig. 4-B). We found 393 ambiguous boxes and 870 invalid boxes in the Reminiscence. We compare PGA with the Passive model that does not have robot-object interaction. Regarding propagation to ambiguous boxes, PGA showed a reduced rate of 41.2% compared with that of 50.4% from the Passive model. Moreover, Passive model had a high tendency to propagate towards invalid boxes, in at 83.7%, while PGA showed nearly negligible rate at 0.07%.
TABLE II: Results of Online Experiment (LCRG) in GraspMine.
Heterogeneous Homogeneous
Method grounding grasping grounding grasping
Direct 63.3 53.3 26.7 20.0
PGA (ours) 90.0 76.7 66.7 53.3
Supervised 100.0 83.3 83.3 70.0
Figure 4: Qualitative Analysis. The top row showcases examples of objects from each phase. In the bottom row, personalized object grounding results by PGA are depicted in blue boxes and those by Passive model in red, set alongside the ground truth in black boxes. Within the Propagation through Reminiscence on the top row, solid lined boxes are the pseudo-labelled objects from the Reminiscence and dotted boxes denote objects that were NOT pseudo-labelled with according indicators in the respective models.
IV-D Online Experiment
Robotic Platform. Our robotic platform is equipped with the 6-DoF Kinova Gen3 Lite [^1] https://www.kinovarobotics.com/product/gen3-lite-robots, complemented by the Intel Realsense Depth Camera D435. While all learning and inferences of PGA are managed on the remote server, manipulation planning is computed on the local platform.
Evaluation Protocol. From the Heterogeneous and Homogeneous splits, we randomly selected 30 images each and closely reproduced the depicted scenes. We assessed LCRG based on two human-evaluated metrics: object grounding success rate and grasping accuracy, i.e., LCRG.
Results. The results, presented in Tab. II, reveal that PGA significantly surpassed the Direct model in grasping, along with enhanced grounding accuracy. The Direct model, which solely relies on object information acquired from human-robot interaction without leveraging raw images from Reminiscence, lagged behind PGA with a performance gap of 23.4 percent points in the Heterogeneous split and 33.3 percent points in the Homogeneous split. Notably, PGA’s performance was comparable to the Supervised method, which necessitates extensive labeling of raw Reminiscence images, highlighting our method’s efficacy with only requiring single human-robot interaction to transfer personal knowledge.
IV-E Qualitative Analysis
Through qualitative analysis, shown in Fig. 4, we visualize examples from the various phases including Object Information Acquisition, Propagation through Reminiscence, and Personalized Object Grounding. For a comprehensive understanding, we compare the performance of PGA against the Passive model, which shows the most strong performance among the baselines. The Passive model does not leverage the data acquired from robot-object interaction.
One striking observation is PGA’s adeptness at inspecting objects from diverse angles as seen in ‘robot-object interaction’. This allows PGA to recognize variations in object orientation, e.g., views from the rear or the base. In instances such as those represented in (c), (d), (e), and (f), the Passive model often overlooks objects in the Reminiscence that differ from their presentations in the human-robot interaction phase. We conjecture that the omitted samples, e.g., (d), (e), and (f), in the training phase affect some wrong prediction results as in C and D. Moreover, as detailed in Sec. IV-C, the Passive model occasionally tends to propagate towards ambiguous boxes in the Reminiscence, as depicted in (a). Such behavior can lead to inaccurate grounding, as observed in A. The tendency of propagating towards incorrect boxes might have affected in some results shown in panel B.
Despite PGA’s robust grounding abilities as seen through A-D, challenges persist in some complex scenarios, such as in a scenario like E, where objects are situated in highly cluttered environments. Nevertheless, our qualitative analysis reveals PGA’s successful Propagation through Reminiscence, enhanced by the robot-object interaction, yielding promising results in Personalized Object Grounding.
V CONCLUSION
This paper introduces a challenging task scenario GraspMine, along with a novel robotic framework PGA to equip robots with the capacity to learn a user’s personalized knowledge through a single human-robot interaction. By expanding the mutual knowledge between a robot and its user beyond the generic, our approach promotes a shift from robot-centric to user-centric interactions. We anticipate that our method will foster more natural, intuitive interaction and significantly enhance the non-expert user’s experience.
Acknowledgements. We extend our gratitude to Jungmin Lee for her invaluable contribution to video editing, and all the reviewers for their insightful comments and feedback.